OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x^2 – 4x – 60 A. –14, 5 B. –5, 14 C. –6, 10 D. –10, 6

OpenStudy (ranga):

x^2 – 4x – 60 = x^2 - 10x + 6x - 60 = x(x-10) + 6(x-10) = (x-10)(x+6) To find the zeros of the polynomial, equate it to zero and solve for x: (x-10)(x+6) = 0 x = 10 or -6. So C is the correct answer.